# Dynamic Speculation Lookahead Accelerates Speculative Decoding of Large Language Models

### v4

@@ Abstract
Speculative decoding is commonly used for reducing the inference latency of large language models. Its effectiveness depends highly on the speculation lookahead (SL)—the number of tokens generated by the draft model at each iteration. In this work we show that the common practice of using the same SL for all iterations (static SL) is suboptimal. We introduce DISCO (DynamIc SpeCulation lookahead Optimization), a novel method for dynamically selecting the SL. Our experiments with four datasets show that DISCO reaches an average speedup of 10% compared to the best static SL baseline, while generating the exact same text. [^1] We will publicly release our code upon publication.
Dynamic Speculation Lookahead Accelerates
Speculative Decoding of Large Language Models
Jonathan Mamou  Oren Pereg  Daniel Korat Moshe BerchanskyNadav TimorMoshe WasserblatRoy Schwartz Intel Labs, Israel Weizmann Institute of Science, Israel School of Computer Science & Engineering, Hebrew University of Jerusalem Correspondence: jonathan.mamou@intel.com
@@ 1 Introduction
Large language models (LLMs) generate tokens autoregressively (Radford et al., 2019; Brown et al., 2020), which often leads to slow generation. Speculative decoding algorithms Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) reduces the inference latency by splitting the inference of a given model into two steps. First, a fast draft model generates tokens autoregressively. Then a more accurate (target) model validates all generated draft tokens simultaneously. See Fig. 1 for an example. The effectiveness of speculative decoding depends on the speculation lookahead (SL)—the number of tokens generated by the draft model at each iteration. An SL too small leads to too many target forwards steps; SL values too large add redundant draft forward passes. Yet, existing speculative decoding approaches use a static SL—an SL that remains constant across all iterations (Leviathan et al., 2023; Miao et al., 2024; Chen et al., 2023; Cai et al., 2024; Li et al., 2024).
This work starts by defining an oracle SL—a method that assigns each iteration its optimal SL. We observe that the optimal SL shows a high variance across iterations (Fig. 2). We then use this oracle to estimate an upper bound of the expected speedup (compared to using static SLs), showing a potential gain of up to 39% speedup. We then propose DISCO, a novel method for selecting the SL before each iteration. DISCO estimates the likelihood of the next draft token being accepted by the target model, and halts the draft model if this likelihood is too small. We evaluate DISCO across various tasks: code generation, text summarization, and instruction following. Our results show an average speedup of 10% compared to optimal static SL and 31% compared to a previously known heuristic for controlling SLs Gante (2023), all without modifying the output text (Appendix F). DISCO also transfers well across tasks from the same category: training it on one task and using it on another leads to similar speedups.
@@ 2 Background: Speculative Decoding
The effectiveness of speculative decoding in accelerating the token generation process relies heavily on the SL parameter, which determines how many tokens are generated by the draft model before each validation step. The effect of the SL on the overall speedup is subject to a tradeoff; higher SLs potentially reduce the number of target model validations, but also increase the number of redundant draft generations (Fig. 1), and vice-versa. The majority of speculative decoding approaches Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023); Cai et al. (2024); Li et al. (2024) use a static SL—the same number of draft tokens are generated per speculative iteration. [^2] A notable exception is Gante (2023), who applies a heuristic for dynamic SL adaptation by modifying the SL based on the acceptance rate of previous iterations. Chen et al. (2023) explored various static SLs, across different target-draft model pairs and tasks, and empirically showed that as the SL rises, the overall speedup increases until reaching a certain threshold, beyond which it either levels off or even regresses. To study the effect of the SL, Leviathan et al. (2023) defined the improvement factor (IF) as the expected token generation speedup:
@@ 3 Dynamic Speculation Lookahead
We start by employing an oracle for detecting the optimal value of SL ($\gamma$) for each speculative iteration. The oracle uses the draft model to autoregressively generate tokens until a mismatch occurs between the predicted tokens of the draft and target models. This process is repeated for each speculative iteration, ultimately returning the optimal (maximum) number of accepted draft tokens per iteration. The mismatch between the tokens is determined by using the rejection sampling algorithm introduced by Leviathan et al. (2023) with zero temperature. This oracle fulfills the speculative decoding potential: generating the maximal number of valid draft tokens at each iteration, and making a minimal number of calls to both draft and target model.
Figure 2 shows the oracle SL values across the speculative iterations for one MBPP example. Compared to the static SL, we observe a lower number of both draft and target forward passes. Figure 3 shows the average oracle SL over the speculative iterations for the Alpaca dataset Taori et al. (2023). Both figures show a high variance of oracle SL values, implying that astatic SL is likely to be suboptimal. See App. A for further analysis.
@@
$C_{i}=FFN(Concat(Top_{k}(y^{D}_{i}),Ent(y^{D}_{i}),i))$
where $Top_{k}()$ selects the top $k$ values and $Ent()$ is the entropy function. [^4] We use $k=10$ in all experiments.
@@ 4 Experiments
Datasets
We evaluate our method on four datasets spanning three tasks: code generation using MBPP Austin et al. (2021) and HumanEval Chen et al. (2021); text summarization using CNN-DailyMail Nallapati et al. (2016); and instruction-following using Alpaca. We use the training sets for training the SL classifier and the validation sets for setting the threshold $\tau$ and the $\text{SL}_{\max}$ hyperparameter. For HumanEval, which has no training and validation sets, we evaluate transfer learning from MBPP. See App. B for more details on the datasets and prompts.
Models
For code generation tasks, we use the Starcoder model family Li et al. (2023)—15B for target and 168M for draft. For the other tasks, we use Vicuna models—13B as target Chiang et al. (2023) and 68M as draft Yang et al. (2024). See App. C for more details.
@@
Baselines
We compare the LLM inference latency of DISCO to both static SL (static SL-5) and dynamic heuristic SL setups (dynHeur SL; Gante, 2023). We also consider the optimal static SL baseline tuned on our validation sets (static SL-opt). Finally, we also report results for our oracle (Sec. 2), which represents the lower bound on latency.
Results
Benchmark Method Latency(ms) Speedup
@@ Results
Table 1: Average latency results on different benchmarks. HumanEval results use a classifier trained on MBPP (transfer learning). All results are provided with greedy decoding (temperature=0).
Table 1 presents our results using the rejection sampling scheme with greedy decoding (temperature=0) since baselines get higher speedup Leviathan et al. (2023). Employing an SL classifier consistently outperforms all other baselines across all benchmarks. Average latency improvements of DISCO over the optimal static SL and the dynamic heuristic baselines are 10.3% and 31.4% respectively, while preserving the same output as the target model. Importantly, our improvement does not come only from our training data: the optimal static SL (as fit by that data) is still underperformed by DISCO. Finally, DISCO transfers well across tasks: when trained on MBPP, it is still outperforms all baselines on HumanEval.
Analysis
Since DISCO is sampling temperature agnostic, we present in Table 2 latency results for CNN-DM with non-zero temperature; we observe that DISCO speedup also improves in that case. In addition, we report latency results using a simple rule-based approach: we use the perplexity to measure the confidence in the sequence of tokens predicted by the draft, when lower perplexity indicates higher confidence. Optimal perplexity threshold is tuned on our validation sets (ppl SL-opt). We observe that it yields a lower speedup.
Benchmark Method Latency(ms) Speedup
CNN-DM Target 36.32 1.00x
static SL-5 21.88 1.66x
DISCO (ours) 19.96 1.82x
ppl SL-opt 26.79 1.43x
@@ 5 Related Work
Pioneering studies on speculative decoding Leviathan et al. (2023); Chen et al. (2023) introduced a rejection sampling scheme that preserves the distribution of the target model, guaranteeing that speculative decoding maintains the quality of the target model. Subsequent work Miao et al. (2024) elevated the average number of accepted tokens by using several draft models. Most recently, Timor et al. (2024) introduced DSI, a distributed variation of speculative decoding that is provably faster than non-distributed methods and does not require additional training or architectural changes. DSI orchestrates multiple servers, and the maximum number of servers is inversely proportional to the SL. To eliminate the need for a separate draft model, Li et al. (2024); Cai et al. (2024); Bhendawade et al. (2024); Yang et al. (2024) train additional, specialized draft layers on top of the transformer decoder. DISCO transfers well within domains and does not require classifier training per dataset whereas these methods necessitate training per dataset. At inference time, they employ a static SL; we believe that DISCO can be beneficially applied to these approaches, we leave this research for future work. Zhang et al. (2023) proposed draft-exiting with an adaptive threshold for self-speculative decoding using a rule-based approach that compares a confidence to a predetermined threshold. This method seems suitable to approaches where the draft is a subset of the target model, whereas our approach is more generic. A very recent concurrent work by S et al. (2024) enhanced the draft model’s accuracy by granting it access to the target model’s representations. In addition, it employed a classifier to determine whether to halt or continue the speculation process. Our work delves into the impact of the SL on the efficiency of speculative decoding, encompassing comparisons between static and dynamic SL approaches, as well as the upper bound of improvement represented by the oracle SL.
@@ 6 Conclusion
7 Limitations
While our study shows improved effectiveness of speculative decoding dynamic SL adaptation, several limitations should be acknowledged.
First, although we conducted experiments on four different datasets, our evaluation was limited to only two pairs of models, namely Vicuna and Starcoder model families. Extending the analysis to a broader range of model configurations could provide a more comprehensive understanding of the generalizability of our findings.
Second, we note that the effectiveness of dynamic speculation lookahead (SL) approaches is contingent upon the standard deviation of the SL as measured by the oracle. When the standard deviation is low, there may be less room for improvement using dynamic SL strategies.
8 Potential Risks
As this work is dedicated to evaluating existing models on publicly available datasets, we are not aware of any potential negative impact.
@@ Appendix A Oracle SL Analysis
Table 3 shows the oracle SL expectancy and standard deviation of the oracle measured on different datasets and models and Figure 4 shows the probability distribution of the oracle SL for the different datasets. We observe a high variance of SL values.
@@ B.1 MBPP
Concerning the prompt, we followed Ben Allal et al. (2022); Fried et al. (2023) and included the description of the programming task and a single test to verify solution, in order to help the model catch the signature of the function (see Fig. 6).
@@ B.4 Alpaca
We follow Taori et al. (2023) to define the prompts. For samples with a non-empty input field, we use the prompt as in Fig. 8 while for samples with empty input field, we use the prompt as in Fig. 9.
@@ D.2 Classifier F1 Results
We evaluate the quality of the classifier by measuring its F1 score on the validation set and report in Table 5 F1 scores for both static SL and DISCO. F1 scores of DISCO always outperforms F1 scores of static SL.

### v5

@@ Abstract
Speculative decoding is commonly used for reducing the inference latency of large language models. Its effectiveness depends highly on the speculation lookahead (SL)—the number of tokens generated by the draft model at each iteration. In this work we show that the common practice of using the same SL for all iterations (static SL) is suboptimal. We introduce DISCO (DynamIc SpeCulation lookahead Optimization), a novel method for dynamically selecting the SL. Our experiments with four datasets show that DISCO reaches an average speedup of 10% compared to the best static SL baseline, while generating the exact same text.
@@ 1 Introduction
Large language models (LLMs) generate tokens autoregressively (Radford et al., 2019; Brown et al., 2020), which often leads to slow generation. Speculative decoding algorithms Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) reduces the inference latency by splitting the inference of a given model into two steps. First, a fast draft model generates tokens autoregressively. Then a more accurate (target) model validates all generated draft tokens simultaneously. See Fig. 1 for an example. The effectiveness of speculative decoding depends on the speculation lookahead (SL)—the number of tokens generated by the draft model at each iteration. An SL too small leads to too many target forwards steps; SL values too large add redundant draft forward passes. Yet, existing speculative decoding approaches use a static SL—an SL that remains constant across all iterations (Leviathan et al., 2023; Miao et al., 2024; Chen et al., 2023; Cai et al., 2024; Li et al., 2024). This work starts by defining an oracle SL—a method that assigns each iteration its optimal SL. We observe that the optimal SL shows a high variance across iterations (Fig. 3). We then use this oracle to estimate an upper bound of the expected speedup (compared to using static SLs), showing a potential gain of up to 39% speedup. We then propose DISCO, a novel method for selecting the SL before each iteration. DISCO estimates the likelihood of the next draft token being accepted by the target model, and halts the draft model if this likelihood is too small. We evaluate DISCO across various tasks: code generation, text summarization, and instruction following. Our results show an average speedup of 10% compared to optimal static SL and 31% compared to a previously known heuristic for controlling SLs Gante (2023), all without modifying the output text (App. F). DISCO also transfers well across tasks from the same category: training it on one task and using it on another leads to similar speedups.
@@ 2 Background: Speculative Decoding
The effectiveness of speculative decoding in accelerating the token generation process relies heavily on the SL parameter, which determines how many tokens are generated by the draft model before each validation step. The effect of the SL on the overall speedup is subject to a tradeoff; higher SLs potentially reduce the number of target model validations, but also increase the number of redundant draft generations (Fig. 1), and vice-versa. The majority of speculative decoding approaches use a static SL—the same number of draft tokens are generated per speculative iteration. [^1] A notable exception is Gante (2023), who applies a heuristic for dynamic SL adaptation by modifying the SL based on the acceptance rate of previous iterations. Chen et al. (2023) explored various static SLs, across different target-draft model pairs and tasks, and empirically showed that as the SL rises, the overall speedup increases until reaching a certain threshold, beyond which it either levels off or even regresses. To study the effect of the SL, Leviathan et al. (2023) defined the improvement factor (IF) as the expected token generation speedup:
@@ 3 Dynamic Speculation Lookahead
We start by employing an oracle for detecting the optimal value of SL ($\gamma$) for each speculative iteration. The oracle uses the draft model to autoregressively generate tokens until a mismatch occurs between the predicted tokens of the draft and target models. This process is repeated for each speculative iteration, ultimately returning the optimal (maximum) number of accepted draft tokens per iteration. The mismatch between the tokens is determined by using the rejection sampling algorithm introduced by Leviathan et al. (2023) with zero temperature. This oracle fulfills the speculative decoding potential: generating the maximal number of valid draft tokens at each iteration, and making a minimal number of calls to both draft and target model. Figure 3 shows the oracle SL values across the speculative iterations for one MBPP example. Compared to the static SL, we observe a lower number of both draft and target forward passes. Figure 3 shows the average oracle SL over the speculative iterations for the Alpaca dataset Taori et al. (2023). Both figures show a high variance of oracle SL values, implying that astatic SL is likely to be suboptimal. See App. A for further analysis.
@@
$C_{i}=FFN(Concat(Top_{k}(y^{D}_{i}),Ent(y^{D}_{i}),i))$ (2)
@@ 4 Experiments
Datasets and Models
We evaluate our method on four datasets spanning three tasks: code generation using MBPP Austin et al. (2021) and HumanEval Chen et al. (2021); text summarization using CNN-DailyMail Nallapati et al. (2016); and instruction-following using Alpaca. We use the training sets for training the SL classifier and the validation sets for setting the threshold $\tau$ and the $\text{SL}_{\max}$ hyperparameter. For HumanEval, which has no training and validation sets, we evaluate transfer learning from MBPP. For code generation tasks, we use the Starcoder model family Li et al. (2023)—15B for target and 168M for draft. For the other tasks, we use Vicuna models—13B as target Chiang et al. (2023) and 68M as draft Yang et al. (2024). See Apps. B and C for more details.
@@
Baselines and Results
We compare the LLM inference latency of DISCO to both static SL (static SL-5) and dynamic heuristic SL setups (dynHeur SL; Gante, 2023).
Benchmark Method Latency Speedup
@@
Table 1: Average latency results (in milliseconds) on different benchmarks. HumanEval results use a classifier trained on MBPP (transfer learning). All results are provided with greedy decoding (temperature=0).
We also consider the optimal static SL baseline tuned on our validation sets (static SL-opt). Finally, we also report results for our oracle (Sec. 2), which represents the lower bound on latency. Table 1 presents our results using the rejection sampling scheme with greedy decoding (temperature=0) since baselines get higher speedup Leviathan et al. (2023). Employing an SL classifier consistently outperforms all other baselines across all benchmarks. Average latency improvements of DISCO over the optimal static SL and the dynamic heuristic baselines are 10.3% and 31.4% respectively, while preserving the same output as the target model. Importantly, our improvement does not come only from our training data: the optimal static SL (as fit by that data) is still underperformed by DISCO. Finally, DISCO transfers well across tasks: when trained on MBPP, it is still outperforms all baselines on HumanEval. See App. G for further analysis.
@@ 5 Related Work
Pioneering studies on speculative decoding Leviathan et al. (2023); Chen et al. (2023) introduced a rejection sampling scheme that preserves the distribution of the target model, guaranteeing that speculative decoding maintains the quality of the target model. Subsequent work Miao et al. (2024) elevated the average number of accepted tokens by using several draft models. Most recently, Timor et al. (2024) introduced DSI, a distributed variation of speculative decoding that is provably faster than non-distributed methods and does not require additional training or architectural changes. To eliminate the need for a separate draft model, Li et al. (2024); Cai et al. (2024); Bhendawade et al. (2024); Yang et al. (2024) train additional, specialized draft layers on top of the transformer decoder. DISCO transfers well within domains and does not require classifier training per dataset whereas these methods necessitate training per dataset. At inference time, they employ a static SL; we believe that DISCO can be beneficially applied to these approaches, we leave this research for future work. Zhang et al. (2023) proposed draft-exiting with an adaptive threshold for self-speculative decoding using a rule-based approach that compares a confidence to a predetermined threshold. This method seems suitable to approaches where the draft is a subset of the target model, whereas our approach is more generic. A very recent concurrent work by S et al. (2024) enhanced the draft model’s accuracy by granting it access to the target model’s representations. In addition, it employed a classifier to determine whether to halt or continue the speculation process. Our work delves into the impact of the SL on the efficiency of speculative decoding, encompassing comparisons between static and dynamic SL approaches, as well as the upper bound of improvement represented by the oracle SL.
@@ Appendix A Oracle SL Analysis
Table 2 shows the oracle SL expectancy and standard deviation of the oracle measured on different datasets and models and Fig. 4 shows the probability distribution of the oracle SL for the different datasets. We observe a high variance of SL values.
@@ B.1 MBPP
Concerning the prompt, we followed Ben Allal et al. [2022], Fried et al. [2023] and included the description of the programming task and a single test to verify solution, in order to help the model catch the signature of the function (see Fig. 6).
@@ B.4 Alpaca
We follow Taori et al. [2023] to define the prompts. For samples with a non-empty input field, we use the prompt as in Fig. 8 while for samples with empty input field, we use the prompt as in Fig. 9.
@@ D.2 Classifier F1 Results
We evaluate the quality of the classifier by measuring its F1 score on the validation set and report in Tab. 4 F1 scores for both static SL and DISCO. F1 scores of DISCO always outperforms F1 scores of static SL.
F1 score measures the accuracy of the classifier in predicting the speculative length (SL) but does not account for how well the predictions align with the oracle’s behavior in reducing latency. In particular, F1 is influenced by 2 different types of error FP and FN that have a different impact on the speedup. FP Errors lead to unnecessary speculative execution, which wastes resources but might not drastically reduce overall speedup. FN Errors lead to missed opportunities for speedup, having a more severe impact on overall latency reduction.
@@ Appendix F Additional Results
Appendix G Further Latency Results Analysis
Concerning SL-opt and SL-5 speedup values on CNN-DM reported in Tab. 1, note that the optimal static SL is tuned on the validation set while the latency and speedup reported numbers are on the test set. This is similar to our setup, where the classifier is tuned on the validation set, and similarly reported on the test set. We observe that for CNN-DM, the optimal SL on the validation set is not as good as the SL-5 on the test set.
Since DISCO is sampling temperature agnostic, we present in Tab. 6 latency results for CNN-DM with non-zero temperature; we observe that DISCO speedup also improves in that case. In addition, we report latency results using a simple rule-based approach: we use the perplexity to measure the confidence in the sequence of tokens predicted by the draft, when lower perplexity indicates higher confidence. Optimal perplexity threshold is tuned on our validation sets (ppl SL-opt). We observe that it yields a lower speedup.
Benchmark Method Latency(ms) Speedup
CNN-DM Target 36.32 1.00x
static SL-5 21.88 1.66x
DISCO (ours) 19.96 1.82x
ppl SL-opt 26.79 1.43x